The following probability distribution graph has a missing value for the outcomes of x = 2 and x = 6. Note that x = {1, 2, 3, 4, 5, 6}.
Furthermore, P(x = 2) + 3P(x = 6) = 0.7
Determine the probability outcomes of both P(x = 2) and P(x = 6). Show all your work.
GRAPH:

0.4-

0.3- ---

0.2- ---
---
0.1-
1 2 3 4 5 6 2&6 are blank.
1
Expert's answer
2020-05-25T18:33:47-0400
"\begin{matrix}
 x & 1 & 2 & 3 & 4 & 5 & 6 \\
 p(x) & 0.1 & a & 0.2 & 0.3 & 0.1 & b
\end{matrix}"

"0.1+a+0.2+0.3+0.1+b=1""a+3b=0.7"

"a+b=0.3""a+3b=0.7"

"a+b=0.3""2b=0.4"

"a=0.1""b=0.2"

"\begin{matrix}
 x & 1 & 2 & 3 & 4 & 5 & 6 \\
 p(x) & 0.1 & 0.1 & 0.2 & 0.3 & 0.1 & 0.2
\end{matrix}"

"P(X=2)=0.1"

"P(X=6)=0.2"